A circle passes through the point (-1,3)&(5,11)and radias is 5 .then its centre is

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Let the center is (h,k)
=> (h+1)2 + (k-3)2 = (h-5)2 +(k-11)2 =52
=> (h+1)2 + (k-3)2 = (h-5)2 +(k-11)2
simplify
3h +4k=34 ...............1
 
(h+1)2 + (k-3)2 =52
h2 +k2 +2h -6k  =15 ..............2
 
Solving equation 1 and equation 2 we get,
h2 + ((34 – 3h)/4)2 +2h – 3(34 – 3h)/2  =15
On solving frther we get,
h = 5 and k = -3